OpenStudy (appleduardo):

what is the derivative of 1/x ?

OpenStudy (anonymous):

1/x = x^(-1) so the derivative is -x^ (-2)

OpenStudy (appleduardo):

why 1 turned negative?

OpenStudy (anonymous):

because that's what you have when you have negative exponents... x^(-1) = 1/x

OpenStudy (anonymous):

^that